OpenStudy (anonymous):

A number is multiplied by 6. If 8 is subtracted from the product you get 34.

OpenStudy (anonymous):

x*6-8=34 What is x?

OpenStudy (anonymous):

Make the number x: 6x-8=34 Can you do that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

7, 34+8=42 42/6=7

OpenStudy (anonymous):

OrangeMaster, the purpose of this website is not to spoonfeed them answers.

OpenStudy (anonymous):

Simple, just reverse the problem to get the answer. The answer is 7

OpenStudy (anonymous):

I didn't, I just told them how to get the answer of a similar problem. Just saying 7 would have been me just giving the answers.

OpenStudy (anonymous):

thanks I got it, it was much easier to reverse it

OpenStudy (anonymous):

Please read http://openstudy.com/study#/updates/527023d6e4b08ad7c0d7b5bc

OpenStudy (anonymous):

6n - 8= 34 6n = 42 n= 7

OpenStudy (anonymous):

@zimmah, I took one look. By the way, I'm SO not reading that. My method worked just as well, and was three times faster.

OpenStudy (anonymous):

I'm not saying that your way was wrong, but it just wasn't my way.

OpenStudy (anonymous):

Your method may have answered the question, but that's not the point of this website. This is not yahoo answers, this is a website designed for educational purposes. Just giving the answer is NOT a good way of teaching someone.

OpenStudy (anonymous):

In fact, you even break the code of conduct by doing that.

OpenStudy (anonymous):

Then please just go to yahoo answers if that is your way, as you obviously don't get the point of this website.

OpenStudy (anonymous):

I did NOT give the answer. I worked it out and told him how he could answer similar problems in the future. I tried being nice, but I'm just going to leave this post because you;re annoying me.

OpenStudy (anonymous):

You did give the answer

OpenStudy (anonymous):

*you're, not you;re

OpenStudy (anonymous):

nope, I didn't

OpenStudy (anonymous):

If you think that I did something so wrong then report me.

OpenStudy (anonymous):

tag a moderator if you want, it's fine with me.

OpenStudy (anonymous):

Are you finished talking now? @zimmah

OpenStudy (anonymous):

@thomaster